the ratio of A to B is 1.37 to 1 and there are 53.2g of A. How many grams of B are predicted by this ratio?
Answer by richwmiller(17219)

a/b=1.37/1,
a=53.2
53.2/b=1.37/1
53.2/b=1.37
53.2=1.37b
53.2/1.37=b